Due to erratic sales of its sole product—a high-capacity battery for laptop computers—PEM, Inc., has been experiencing financial difficulty for some time. The company’s contribution format income statement for the most recent month is given below.

Sales (12,700 units × $20 per unit)

$

254,000

Variable expenses

152,400

Contribution margin

101,600

Fixed expenses

113,600

Net operating loss

$

(12,000

)

Required:

1. Compute the company’s CM ratio and its break-even point in unit sales and dollar sales.

2. The president believes that a $6,200 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in an $84,000 increase in monthly sales. If the president is right, what will be the increase (decrease) in the company’s monthly net operating income?

3. Refer to the original data. The sales manager is convinced that a 10% reduction in the selling price, combined with an increase of $38,000 in the monthly advertising budget, will double unit sales. If the sales manager is right, what will be the revised net operating income (loss)?

4. Refer to the original data. The Marketing Department thinks that a fancy new package for the laptop computer battery would grow sales. The new package would increase packaging costs by $0.70 per unit. Assuming no other changes, how many units would have to be sold each month to attain a target profit of $4,500?

5. Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $52,000 each month.

a. Compute the new CM ratio and the new break-even point in unit sales and dollar sales.

b. Assume that the company expects to sell 20,300 units next month. Prepare two contribution format income statements, one assuming that operations are not automated and one assuming that they are. (Show data on a per unit and percentage basis, as well as in total, for each alternative.)

c. Would you recommend that the company automate its operations (Assuming that the company expects to sell 20,300)?
Answer #1

Solution 1:

CM ratio = Contribution margin / Sales = $101,600 / $254,000 = 40%

break-even point in unit sales = Fixed expenses / CM per unit = $113,600 / $8 = 14200 units

Breakeven point in dollar sales = Fixed expenses / CM ratio = $113,600 / 40% = $284,000

Solution 2:

increase (decrease) in the company’s monthly net operating income = Increase in contribution margin - Increase in advertising expenses

= ($84,000*40%) - $6,200 = $27,400

Solution 3:

New selling price per unit = $20*90% = $18

New CM per unit = $8 - $2 = $6 per unit

New fixed costs =$113,600 + $38,000 = $151,600

New sales volume = 12700*2 = 25400 units

New operating income = Contribution margin - Fixed costs = (25400*$6) - $151,600 = $800

Solution 4:

New contribution margin per unit = $8 - $0.70 = $7.30 per unit

Target income = $4,500

Nos of units to be sold to attain target profit = (Fixed cost + Net operating income) / CM per unit

= ($113,600 + $4,500) / $7.30 = 16178 units

Solution 5a:

New CM per unit = $8 +$3 = $11 per unit

New fixed costs = $113,600 + $52,000 = $165,600

New CM ration = $11 / $20 = 55%

New break even point in units = $165,600 / $11 = 15055 units

Breakeven point in dollar sales = $165,600 /55% = $301,091

Solution 5b:

Contribution format income statement - Operation Automated
Particulars Per unit % Total
Sales $20.00 100% $406,000.00
Variable costs $9.00 45% $182,700.00
Contribution margin $11.00 55% $223,300.00
Fixed costs $165,600.00
Net operating income $57,700.00
Contribution format income statement - Operation not Automated
Particulars Per unit % Total
Sales $20.00 100% $406,000.00
Variable costs $12.00 60% $243,600.00
Contribution margin $8.00 40% $162,400.00
Fixed costs $113,600.00
Net operating income $48,800.00

Solution 5c:

Yes, it is recommended company should automate its operations